# Improving Efficiency of Diffusion Models via Multi-Stage Framework and Tailored Multi-Decoder Architectures

### v1

@@ 3.2 Optimal Denoiser-based Timestep Clustering
such that the average functional similarity of $\bm{\epsilon}^{*}_{t}$ (resp. $\bm{\epsilon}^{*}_{t}$) to $\bm{\epsilon}^{*}_{0}$ (resp. $\bm{\epsilon}^{*}_{1}$) in $[0,t_{1})$ (resp. $[t_{2},1]$) is larger than or equal to a pre-defined threshold $\alpha$. As the above optimization problems are intractable, we propose the procedure outlined in Algorithm 1 to obtain an approximate solution. [^3] Basically, the algorithm samples $K$ sample pairs $(\bm{y}_{k},\bm{\epsilon}_{k},t_{k}),k\in\{1,\ldots,K\}$ to calculate the distance $\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{0},\bm{y}_{k},\bm{\epsilon}_{k})$ and $\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{1},\bm{y}_{k},\bm{\epsilon}_{k})$ and then based on those distances to solve the optimization problem define in the line 10 -11 of Algorithm 1 to get $t_{1}$ and $t_{2}$.
4 Identification of Key Sources of Inefficiency
Conventional diffusion model architectures, as exemplified by [2, 1, 18], treat the training of the diffusion model as a unified process across all timesteps. Recent research endeavors, such as [19, 20, 21, 22], have highlighted the benefits of recognizing distinctions between different timesteps and the potential efficiency gains from treating them as separate tasks during the training process. However, our experimental results demonstrate that both unified and separate architectures suffer inefficiency for training diffusion models, where the infficiency comes from (i) overfitting and (ii) gradient dissimilarity.
4.1 Empirical Observations on the Key Sources of Inefficiency
(a) Analysis on interval $[0,t_{1})$
(b) Analysis on interval $[t_{2},1]$
Experiment setup.
In our experiments, we consider three-stage training and divide the time steps into three intervals: $[0,t_{1}),[t_{1},t_{2}),[t_{2},1]$. [^4] Details for interval clustering can be found in Section 5. Let $(\bm{\epsilon}_{\bm{\theta}})^{[a,b]}_{i},0\leq a<b\leq 1$ denote a U-Net architecture with parameter $\bm{\theta}$ trained with $i$ iterations and fed with data pairs $(\bm{x}_{t},t)$, where $t\in[a,b]$. We then trained models using two different strategies: a unified architecture with 108M network parameters for all intervals, i.e., $(\bm{\epsilon}_{\bm{\theta}})^{[0,1]}_{i}$, and separate architectures with varying network parameters (e.g., 47M, 108M, 169M) for each interval; e.g., $(\bm{\epsilon}_{\bm{\theta}})^{[0,t_{1})}_{i}$ for the interval $[0,t_{1})$, etc. It is worth noting that, apart from the differences in network parameters, we utilize the same network architecture (e.g., U-Net) for both the unified and the separate approaches. We assessed the training progress of each model by evaluating image generation quality at different training iterations. Notably, because some of the models are only trained on one interval, we need to provide a ground truth score for the other intervals. In Figure 2, the sampling process is shown above and the experimental results are shown below.
Inefficiency in unified architectures.
From Figure 2, we observe the following:
  • Overparameterization and underfitting emerge simultaneously for unified architectures. In Figure 1(a), we observe that increasing the number of parameters in Interval 0 can improve the image generation quality (as indicated by a lower FID score). In contrast, Figure 1(b) reveals that increasing the number of parameters in Interval 2 has minimal impact on the quality of image generation. This implies that using a unified architecture will result in underfitting in Interval 0 and overparameterization in Interval 2. The current unified architecture’s parameter redundancy leaves significant room for improving its efficiency. To optimize the computational usage, we should allocate more parameters to Interval 0 while allocate fewer parameters to Interval 2.
  • Gradient dissimilarity hinders training for unified architecture. Quantitative results from [20] demonstrate dissimilarity in gradients emerges among different intervals. This can also be observed from our results based upon both Figure 1(a) and Figure 1(b). For the unified and separate architectures using the same number of parameters (108M), separate architecture achieves a significantly lower FID with the same training iterations, implying that dissimilar gradients among intervals may hinder training when using a unified architecture. Here, the only difference between training separate and unified architectures is that the batch gradient for unified architecture is calculated based on all timesteps while the batch gradient for separate architecture is calculated only from a specific interval.
Inefficiency in existing separate architectures.
Although separated architecture [19, 21, 22] better allocates computational resources for each interval, it suffers from overfitting. This can be illustrated based upon training separate architectures (169M) and (108M) in Interval 0 shown in Figure 1(a), where increasing the number of parameters will lead to overfitting when we increase the number of training iterations beyond achieving the best FID. This also happens in Interval 2, when we compare all separate architectures in Figure 1(b). In comparison, the unified networks with 108M parameters are less prone to overfit for both Interval 0 and Interval 2. This suggests that we can reduce overfitting by training shared weights across different intervals together.
4.2 Tackling the Inefficiency via Multistage U-Net Architectures
@@ 5 Experiments
For this section, we start with providing the basic experimental setups. Followed by this, in Section 5.1, we demonstrate the improved image generation quality of our multistage architecture over the state-of-the art models, under comparable training and sampling computations. On the other hand, in Section 5.2, we demonstrate the improved training and sampling efficiency of our methods over existing methods, when compared with similar image generation quality. Finally, in Section 5.3 and Section 5.4, we conduct comprehensive ablation studies on timestep clustering methods and multistage architectures, illustrating the superiority of the choices in our method.
@@
5.1 Comparison of Image Generation Quality
In this subsection, we demonstrate the effectiveness of our approach by comparing the image generation quality (measured by FID) with comparable training and sampling computations (measured by NFE). Specifically, Table 1 presents FID scores to measure the sampling quality, and NFEs to measure the number of sampling iterations required using the CIFAR-10 dataset. Our method is compared to 8 baselines. As observed, our multistage DPM-Solver outperforms DPM-Solver in terms of the reported FID values while requiring similar training iterations (both are $4.5\times 10^{5}$), and model GFLOPs (18.65 for multistage DPM-Solver versus 17.65 for DPM-Solver). A similar observation holds when we compare our multistage EDM and the vanilla EDM, where we reduce FID from 2.05 to 1.96 by using the multi-stage architecture. Remarkably, utilizing only 20 NFE, our Multistage DPM-Solver returns the same FID score as the one reported for the PFGM method, which requires 147 NFEs. These results also highlight the adaptability of our framework to higher-order ODE solvers; see the 8th and last row of Table 1.
@@ 5.1 Comparison of Image Generation Quality
5.2 Comparison of Training & Sampling Efficiency
@@
In this subsection, we further demonstrate the superiority of our method by comparing the training and sampling efficiency under comparable image generation quality. Specifically, in Table 2, we present the number of training iterations, GFLOPs, and total training PFLOPs of our approach, DPM-solver, EDM, and LDM using CIFAR-10 and CelebA datasets. Using the CIFAR-10 dataset, our multistage DPM-solver achieves similar FID scores (2.71 vs 2.73) while requiring nearly half the training iterations when compared to the vanilla DPM solver. For the case of EDM (resp. LDM), our approach returns an FID score of 1.44 (resp. 8.29), requiring $1.4\times 10^{5}$ ($3.2\times 10^{5}$) less iterations when compared to vanilla DPM (resp. LDM). For the cases of DPM and EDM, we can achieve substantial reduction of training iterations, which is demonstrated by a marginal increase in the number of GFLOPs (row 3 vs. row 4 and row 5 vs. row 6). For the LDM case,we also achieve a significant reduction of training iterations, demonstrated by a reduced number of GFLOPs (row 8 vs. row 9). These promising results highlight the significantly improved computational efficiency achieved by using the proposed multistage framework.
@@ 5.3 Comparison of Different Architectures
In Section 4, we highlighted the limitations of both unified and separate diffusion model architectures in terms of training efficiency (see Figure 2). In this part, we further illustrate these limitations through extensive experiments as shown in Table 3. Here, we use the U-Net architecture, trained on the CIFAR-10 dataset, and utilize the DPM-Solver for sampling. For the unified case, we use a single U-Net model with 128 channels. For the separate case, three distinct U-Nets are used by which, for each interval, we use a 128-channel decoder. For improved performance of the separate architecture, we implement two techniques: early stopping (ES) and tailored parameters (TP) to tackle the overfitting and parameter inefficiency discussed in Section 4. Under ES, the criteria is to stop training prior to overfitting. For the tailored parameters, the three U-Nets are configured with 192, 128, and 16 channels decoders for Intervals 0, 1, and 2, respectively.
@@
Our comparison and analysis in Table 3 reveals notable insights of our network design. Comparisons between the 2nd and 3rd rows (and between the 4th and 5th rows) on the separate architectures indicate that early stopping effectively mitigates overfitting and enhances generation quality. When comparing the 2nd and 4th rows (as well as the 3rd and 5th rows) on the separate architectures, we observe that optimizing parameter usage can achieve a significant decrease in FID under comparable GFLOPs (18.65 vs. 17.65). Most importantly, our multistage architecture, as shown in the 6th row, benefits from both unified and separate architectures, achieving the best FID (2.35, compared to 2.73 and 2.52). Comparing the 2nd row and 4th row, the shared encoder not only prevents overfitting but also improves the convergence of the diffusion model as reported by the FID scores.
@@ 5.4 Comparison of Timestep Clustering Methods
As preciously stated in Section 2, various timestep clustering methods are proposed including timestep-based, SNR-based, and gradient-based clustering approaches [20, 21]. In this subsection, we conduct an experiment to demonstrate the superiority of our clustering method compared to previous arts. Specifically, we apply the clustering methods in [20, 21] to partition the interval along with our proposed multistage UNet architecture. The computed intervals are shown in the Table 4. We use the multistage DPM-Solver with these different intervals trained on the CIFAR-10 dataset. As observed, our optimal denoiser-based clustering method achieves the highest FID score, consistently outperforming all other clustering methods.
6 Conclusion
In this study, we significantly enhance the training and sampling efficiency of diffusion models through a novel multi-stage framework. This method divides the time interval into several stages, using a specialized multi-decoder U-net architecture that combines time-specific models with a common encoder for all stages. Based on our practical findings, this multi-stage approach utilizes unique parameters for each timestep, along with shared parameters across all timesteps, grouped according to the most effective denoiser. We conducted thorough numerical experiments with three leading-edge diffusion models, such as large-scale latent diffusion models, which confirmed the effectiveness of our strategy.
In future research, it would be interesting to expand our multi-stage approach beyond the unconditional diffusion models, such as conditional diffusion models and diffusion models for solving inverse problems. Our experiments, as detailed in Section 5.2, demonstrate that training latent diffusion models within our multi-stage framework requires only 30% of the computational effort needed for training standard latent diffusion models on the CelebA dataset. Thus, employing a multi-stage strategy could significantly reduce the computational demands for training large-scale stable diffusion models, such as those described in [8], which typically requires significant computational power.
@@ Appendix A Proof of Proposition 1
A.1 Background
This section presents the proof of Proposition 1 in Section 3. Our proof partially follows [18]. As background, let $p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})$ be the perturbation kernel of the diffusion model, which is a continuous process gradually adding noise from original image $\bm{x}_{0}$ to $\bm{x}_{t}$ across $t\in[0,1]$. Both $s_{t}=s(t)$ and $\sigma_{t}=\sigma(t)$ are used interchangeably to denote scalar functions of $t$ that control the perturbation kernel. It has been shown in [2] that this perturbation kernel is equivalent to a stochastic differential equation $\text{d}\bm{x}=\bm{x}_{t}f(t)\text{d}t+g(t)\text{d}\bm{\omega}_{t}$, where $f(t),g(t)$ are scalar functions of $t$. The relations of $f(t),g(t)$ and $s_{t},\sigma_{t}$ are:
$s_{t}=\text{exp}(\int_{0}^{t}f(\xi)\text{d}\xi),\ \ \text{and}\ \sigma_{t}=\sqrt{\int_{0}^{t}\frac{g^{2}(\xi)}{s^{2}(\xi)}\text{d}\xi}\>.$ (9)
A.2 Proof
@@
The above equation is a convex optimization problem by which the solution can be obtained by setting the gradient of $\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)$ w.r.t $\bm{\epsilon}_{\bm{\theta}}(\bm{x};t)$ to zero. This means
@@
Appendix B Generalized Algorithm 1
2:Output: Timesteps $t_{1}$, $t_{2}$, $\ldots$, $t_{n-1}$
3:$\mathcal{S}\leftarrow\emptyset$
4:for $k\in\{1,\dots,K\}$ do
8:end for
[…]
Appendix C Additional Experiments
@@
C.1 Generation Results
The generation samples displayed in Figure 3, from the CelebA and CIFAR-10 datasets, demonstrate that our Multistage strategy is capable of producing high-quality results. Notably, it achieves this with reduced training and computational requirements in comparison to baseline methods.
C.2 Ablation Study on the Network Parameters
@@
C.3 Ablation Study on the Number of Intervals
In this subsection, we conduct ablation studies on the number of intervals of our multistage architecture. We utilize Multistage DPM-Solver training on the CIFAR-10 dataset. To extend Algorithm 1, we design the procedure outlined in Algorithm 2 (Appendix B). The splitting intervals and results are given in Table 7. We compare the stages from one to five. The parameters designed for those architectures are based on the best three-stage model architecture in Table 5. More specifically, we use linear interpolation. Suppose the partitioning timesteps for the three-stage architecture are $t_{1}$ and $t_{2}$ and parameter numbers for each interval are $p_{1},p_{2}$, and $p_{3}$. The parameter $p$ for a random interval $[t_{a},t_{b}]$ is calculated as:
$p=\frac{1}{L([t_{a},t_{b}])}[\mathcal{L}([t_{a},t_{b}]\cap[0,t_{1}))p_{1}~{}+~{}\mathcal{L}([t_{a},t_{b}]\cap[t_{1},t_{2}))p_{2}~{}+~{}\mathcal{L}([t_{a},t_{b}]\cap[t_{2},1])p_{3}]\>,$ (23)
where $\mathcal{L}$ is an operator that measures the length of an interval. We design the parameters for all architectures in Table 7 based on (23), and split intervals for each stage accordingly. As observed, the three-stage architecture, adopted in the main body of the paper, outperforms all other settings.

### v2

@@ 4.2 Optimal Denoiser-based Timestep Clustering
such that the average functional similarity of $\bm{\epsilon}^{*}_{t}$ (resp. $\bm{\epsilon}^{*}_{t}$) to $\bm{\epsilon}^{*}_{0}$ (resp. $\bm{\epsilon}^{*}_{1}$) in $[0,t_{1})$ (resp. $[t_{2},1]$) is larger than or equal to a pre-defined threshold $\alpha$. As the above optimization problems are intractable, we propose the procedure outlined in algorithm 1 to obtain an approximate solution. In particular, the algorithm samples $K$ pairs $(\bm{y}_{k},\bm{\epsilon}_{k},t_{k}),k\in\{1,\ldots,K\}$ to calculate the distances $\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{0},\bm{y}_{k},\bm{\epsilon}_{k})$ and $\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{1},\bm{y}_{k},\bm{\epsilon}_{k})$ (step 6). Based on those distances, we solve the optimization problems defined in the lines 8 and 9 of algorithm 1 to obtain $t_{1}$ and $t_{2}$.
4.3 Rationales for the proposed architecture
Finally, we summarize the rationales of our proposed architecture based on the empirical observations in section 3 and previous works.
Rationales for the shared encoder. (i) Prevent overfitting: If we treat training separate stages of diffusion models as multi-task learning, [35] suggests that shared parameters across tasks in multi-task learning can mitigate overfitting. (ii) Maintain consistency in h-space. The output of the encoder for UNet is named as h-space [36], which has properties for semantic manipulation such as homogeneity, linearity, robustness, and consistency across timesteps. Therefore, sharing the encoder can maintain better consistency of the h-space among all timesteps compared to separate encoders.
Rationales for network parameter design. First, we provide some intuitions as to why the learning task at $t=0$ is harder than that at $t=1$. Suppose diffusion models could converge to the optimal denoiser $\bm{\epsilon}_{\bm{\theta}}^{*}(\bm{x}_{t},t)$ given in eq. 4. Based on this, we observe: (i) when $t\to 0$, we have $\bm{x}_{t\to 0}=\bm{x}_{0}$, $\bm{\epsilon}_{\bm{\theta}}^{*}|_{t\to 0}=\bm{\epsilon}$, so that $\bm{\epsilon}^{*}_{\bm{\theta}}$ is a complicated mapping from training data distribution $p_{\text{data}}(\bm{x})$ to the gaussian distribution; (ii) when $t\to 1$, we have $\bm{x}_{t\to 1}=\bm{\epsilon}$, $\bm{\epsilon}_{\bm{\theta}}^{*}|_{t\to 1}=\bm{\epsilon}$, so $\bm{\epsilon}^{*}_{\bm{\theta}}$ is an identity mapping. The two extreme cases reveal that the identity mapping for $t\to 1$ (close to noise) is easier to learn than that for $t\to 0$. Second, our choice of network parameters across different stages share similar spirits with the recent work [37], which employs a high dimensional subspace as $t\to 0$ and gradually decreases the dimensionality of each subspace until $t=1$.
@@ 5 Experiments
For this section, we start with providing the basic experimental setups. Followed by this, in section 5.1, we demonstrate the improved image generation quality of our multistage architecture over the state-of-the art models, under comparable training and sampling computations. On the other hand, in 5.2, we demonstrate the improved training and sampling efficiency of our methods over existing methods, when compared with similar image generation quality. Finally, in 5.3 and 5.4, we conduct comprehensive ablation studies on timestep clustering methods and multistage architectures, illustrating the superiority of the choices in our method.
@@
5.1 Image Generation Quality Results
In this subsection, we demonstrate the effectiveness of our approach by comparing the image generation quality (measured by FID) with comparable training and sampling computations (measured by NFE). Specifically, table 1 presents FID scores to measure the sampling quality, and NFE to measure the number of sampling iterations required using the CIFAR-10 dataset. Our method is compared to 8 baselines. As observed, our multistage DPM-Solver outperforms DPM-Solver in terms of the reported FID values while requiring similar training iterations (both are $4.5\times 10^{5}$) and model GFLOPs (18.65 for multistage DPM-Solver versus 17.65 for DPM-Solver). A similar observation holds when we compare our multistage EDM and the vanilla EDM, where we reduce FID from 2.05 to 1.96 by using the multi-stage architecture. Remarkably, utilizing only 20 NFE, our Multistage DPM-Solver returns the same FID score as the one reported for the PFGM method, which requires 147 NFEs. These results also highlight the adaptability of our framework to higher-order ODE solvers; see the 8th and last row of table 1.
@@ 5.1 Image Generation Quality Results
5.2 Training & Sampling Efficiency Results
@@
In this subsection, we further demonstrate the superiority of our method by comparing the training and sampling efficiency under comparable image generation quality. Specifically, in table 2, we present the number of training iterations, GFLOPs, and total training PFLOPs of our approach, DPM-solver, EDM, and LDM using CIFAR-10 and CelebA datasets. Using the CIFAR-10 dataset, our multistage DPM-solver achieves similar FID scores (2.71 vs 2.73) while requiring nearly half the training iterations when compared to the vanilla DPM-solver. For the case of EDM (resp. LDM), our approach returns an FID score of 1.44 (resp. 8.29), requiring $1.4\times 10^{5}$ ($3.2\times 10^{5}$) less iterations when compared to vanilla DPM-solver (resp. LDM). For the cases of DPM-solver and EDM, we can achieve a substantial reduction of training iterations, which is demonstrated by a marginal increase in the number of GFLOPs. For the LDM case, we also achieve a significant reduction of both training iterations and GFLOPs. These promising results highlight the significantly improved computational efficiency achieved by using the proposed multistage framework.
@@ 5.3 Comparison of Different Architectures
In section 3, we highlighted the limitations of both unified and separate diffusion model architectures in terms of training efficiency (see fig. 2). In this part, we further illustrate these limitations through extensive experiments as shown in table 3. Here, we use the U-Net architecture, trained on the CIFAR-10 dataset, and utilize the DPM-Solver for sampling. For the unified case, we use a single U-Net model with 128 channels. For the separate case, three distinct U-Nets with 128 channels are used. For improved performance of the separate architecture, we implement two techniques: early stopping (ES) and tailored parameters (TP) to tackle the overfitting and parameter inefficiency discussed in section 3. Under ES, the criteria is to stop training prior to overfitting. For TP, the three U-Nets are configured with 192, 128, and 16 channels decoders for Intervals 0, 1, and 2, respectively.
@@
Our comparison and analysis in table 3 reveal notable insights of our network design. Comparisons between the 2nd and 3rd rows (and between the 4th and 5th rows) on the separate architectures indicate that early stopping effectively mitigates overfitting and enhances generation quality. When comparing the 2nd and 4th rows (as well as the 3rd and 5th rows) on the separate architectures, we observe that optimizing parameter usage can achieve a significant decrease in FID under comparable GFLOPs. Most importantly, our multistage architecture, as shown in the 6th row, benefits from both unified and separate architectures, achieving the best FID (2.35, compared to 2.73 and 2.52). Comparing the 2nd row and the 4th row, the shared encoder not only prevents overfitting but also improves the convergence of the diffusion model as reported by the FID scores.
@@ 5.4 Comparison of Timestep Clustering Methods
As previously stated in section 2, various timestep clustering methods are proposed including timestep-based, SNR-based, and gradient-based clustering approaches [20, 21]. In this subsection, we conduct an experiment to demonstrate the superiority of our clustering method compared to previous arts. Specifically, we apply the clustering methods in [20, 21] to partition the interval along with our proposed multistage UNet architecture. The computed intervals are shown in the table 4. We use the multistage DPM-Solver with these different intervals trained on the CIFAR-10 dataset. As observed, our optimal denoiser-based clustering method achieves the highest FID score, consistently outperforming all other clustering methods.
6 Conclusion & Future Work
In this paper, we introduced a novel multi-stage framework for diffusion models (DM) to improve the training and sampling efficiency. We proposed an algorithm that divides the timestep into several stages. Based on these stages, we designed a stage-specific multi-decoder U-net architecture and a shared encoder across all stages. We conducted thorough numerical experiments with several SOTA diffusion model frameworks and confirmed the effectiveness of our strategy using small scale and large scale datasets.
In future research, it would be interesting to expand our multi-stage approach beyond unconditional diffusion models by considering conditional DMs and DM-based inverse problems solvers. Our experiment in section 5.2 demonstrate that training latent diffusion models within our multi-stage framework requires only 30% of the computational effort needed for training standard latent diffusion models on the CelebA dataset. Thus, employing a multi-stage strategy could significantly reduce the computational demands for training large-scale stable diffusion models, such as those described in [8], which typically requires significant computations.
@@ Appendix A Proof of Proposition 1
This section presents the proof of Prop. 1 in section 4.
@@
The above equation is a convex optimization problem with respect to $\bm{\epsilon}_{\bm{\theta}}$ by which the solution can be obtained by setting the gradient of $\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};\bm{x},t)$ w.r.t $\bm{\epsilon}_{\bm{\theta}}(\bm{x};t)$ to zero. Under the assumption that $\bm{\epsilon}_{\bm{\theta}}$ has an infinite model capacity, and can approximate any continuous function to an arbitrary level of accuracy based on the Universal Approximation Theorem [41], then
@@
Appendix B Additional Experiments
@@
B.1 Samples of Additional Generated Results
The generation samples displayed in fig. 3, from the CelebA and CIFAR-10 datasets, demonstrate that our Multistage strategy is capable of producing high-quality results. Notably, it achieves this with reduced training and computational requirements in comparison to baseline methods.
B.2 Comparison with multi-architecture multi-expert (MEME).
In this subsection, we compare our multistage framework with MEME diffusion model [21]. We reimplement the MEME architecture following their released architecture in Table 4 of [21]. We train both our multistage framework and their MEME on the CIFAR-10 dataset and use DPM-Solver for sampling. The results are shown in table 5. From our experiments, Multistage outperforms MEME in FID with much fewer training iterations and total training PFLOPs. However, because [21] only implements the architecture for CelebA-HQ and FFHQ datasets, the method could potentially achieve better results than our reported results with more hyperparameter tunning. Also, the training is unstable and the loss explodes when we try to increase the model capacity of MEME to be comparable to our Multistage.
Table 5: Training and Sampling Efficiency with related works.
CIFAR-10 $32\times 32$
MEME DPM-Solver [21] $6.0\times 10^{5}$ 13.08 7.85 3.23
Multistage DPM-Solver (Ours) $\textbf{2.5}\times\textbf{10}^{5}$ 18.65 4.66 2.71
B.3 Results in terms of the Precision and Recall Metrics
We evaluate the precision and recall [42] of our proposed architecture on the CIFAR-10 dataset, as results shown in table 6. The recall of the multistage diffusion models is slightly better than their vanilla counterparts. This is potentially because the multistage architecture introduces more degree of freedom when $t=0$ (a separate decoder for this interval), resulting in more diverse data generation.
Table 6: Precision and Recall on CIFAR-10.
METHOD Precision($\uparrow$) Recall($\uparrow$)
DPM-Solver 0.660 0.613
EDM 0.677 0.620
Multistage DPM-Solver 0.657 0.615
Multistage EDM 0.672 0.629
B.4 Ablation Study on the Network Parameters
@@
B.5 Ablation Study on the Number of Intervals
In this subsection, we conduct ablation study on the number of intervals of our multistage architecture. We utilize Multistage DPM-Solver training on the CIFAR-10 dataset. To extend algorithm 1, we propose the procedure outlined in algorithm 2 (appendix C). The splitting intervals and Results are given in table 9. We compare the stages from one to five.